# What Matters to Enhance Traffic Rule Compliance of Imitation Learning for Automated Driving

### v2

@@ IV-E Ablation Studies
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
No CSG $45.67\pm 5.51$ $84.82\pm 2.25$ $0.52\pm 0.10$ $0.02\pm 0.01$ $0.14\pm 0.03$ $0.03\pm 0.02$ $0.09\pm 0.02$ $0.01\pm 0.01$
No Penalty $37.32\pm 3.58$ $78.41\pm 11.32$ $0.54\pm 0.21$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.04\pm 0.02$
$\lambda_{1}=0.7$ $54.36\pm 1.78$ $84.37\pm 5.01$ $0.64\pm 0.04$ $0.01\pm 0.01$ $0.18\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.01$ $0.01\pm 0.00$
$\lambda_{1}=0.3$ $56.80\pm 2.22$ $89.60\pm 1.90$ $0.60\pm 0.02$ $0.01\pm 0.01$ $0.18\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
$\lambda_{2}=0.07$ $51.54\pm 1.50$ $81.49\pm 1.42$ $0.63\pm 0.02$ $0.02\pm 0.01$ $0.13\pm 0.10$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
$\lambda_{2}=0.03$ $54.59\pm 2.88$ $88.12\pm 1.95$ $0.64\pm 0.02$ $0.01\pm 0.01$ $0.11\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.06$ $0.01\pm 0.01$
$\lambda_{3}=0.7$ $55.98\pm 3.00$ $\textbf{89.81}\pm 3.33$ $0.63\pm 0.02$ $0.01\pm 0.01$ $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0\textbf{0.00}\pm 0$ 0.00
$\lambda_{3}=0.3$ $51.98\pm 1.93$ $84.66\pm 3.53$ $0.63\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.08\pm 0.03$ $0.02\pm 0.00$ $0.10\pm 0.02$ $0.01\pm 0.01$
TABLE III: Ablation Study - Town 05 Long Benchmark The ablation study focuses on significant hyper-parameters, namely $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, which correspond to the penalty weights associated with the red light, speed, and stop sign, respectively. Each model is evaluated three times with random seeds. Our optimal model employs default weights of $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, and $\lambda_{3}=0.5$. During the ablation study, we isolate the impact of a single hyper-parameter by modifying it while keeping the others constant.
V Robustness Study
Attack Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
@@ V Robustness Study
FGSM Attack $(\epsilon=0.01)$ P-CSG(Ours) $\mathbf{12.80}\pm 1.35$ $21.70\pm 2.87$ $\mathbf{0.59}\pm 0.02$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.22}\pm 0.23$ $2.40\pm 1.00$ $\mathbf{0.07}\pm 0.11$ $\mathbf{0.00}\pm 0.00$
Transfuser+ $2.69\pm 0.73$ $42.44\pm 3.65$ $0.22\pm 0.06$ $\mathbf{0.00}\pm 0.00$ $1.23\pm 0.31$ $0.26\pm 0.17$ $0.08\pm 0.03$ $0.04\pm 0.02$
Interfuser* $8.37\pm 1.06$ $\mathbf{72.04}\pm 4.15$ $0.15\pm 0.04$ $0.00\pm 0.01$ $0.64\pm 0.27$ $\mathbf{0.10}\pm 0.03$ $0.22\pm 0.03$ $0.07\pm 0.01$
Dot Attack P-CSG(Ours) $\mathbf{36.60}\pm 3.17$ $\mathbf{75.53}\pm 6.43$ $\mathbf{0.56}\pm 0.07$ $0.01\pm 0.02$ $\mathbf{0.06}\pm 0.03$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.10}\pm 0.04$ $\mathbf{0.01}\pm 0.00$
Transfuser+ $0.53\pm 0.21$ $54.06\pm 5.48$ $0.02\pm 0.01$ $\mathbf{0.00}\pm 0.00$ $2.53\pm 0.72$ $0.99\pm 0.16$ $0.21\pm 0.04$ $0.04\pm 0.01$
Interfuser* $12.56\pm 2.15$ $73.62\pm 7.32$ $0.19\pm 0.02$ $0.01\pm 0.01$ $0.68\pm 0.26$ $\mathbf{0.00}\pm 0.00$ $0.21\pm 0.02$ $0.09\pm 0.02$
TABLE IV: Attack Results. This table presents a comparative analysis of the performance and robustness of autonomous driving models under two types of adversarial attacks: FGSM [28] and Dot Attacks [29]. Driving Score serves as the primary metric of overall performance, while the infraction score aggregates various safety-related factors. P-CSG, our proposed model, consistently outperforms the other models across these metrics. Interfuser* has the same PID Controller as P-CSG and Transfuser+ to ensure a fair comparison.
Given the life-critical nature of autonomous driving, it becomes essential to examine the robustness of neural network models against such vulnerabilities.
In this regard, we conduct a comprehensive comparative analysis to evaluate the robustness of our model. We benchmark its performance against two state-of-the-art architectures—Transfuser+ [7] and Interfuser [8]. We focus on two specific types of white-box sensor attacks. These are Fast Gradient Sign Method (FGSM) Attack [28] and Dot Attack [29].
FGSM Attack [28]: FGSM Attack is notable for their computational efficiency and minimal perceptibility to the human eye. This approach exploits the gradient of the model’s loss function with respect to the input data.
$\bm{\eta}=\epsilon_{\textrm{p}}\operatorname{sign}(\nabla_{I}J(\bm{\Theta},\bm{I},y)).$ (15)
$\bm{\tilde{I}}=\bm{I}+\bm{\eta}$ (16)
where $\bm{I}$ is the original image input, and $y$ is the ground true label. The loss function is $J(\bm{\Theta},\bm{I},y)$, and its gradient with respect to ${\bm{I}}$ is $\nabla_{\bm{I}}J(\bm{\Theta},\bm{I},y)$. The hyperparameter $\epsilon_{\textrm{p}}$ controls perturbation magnitude, leading to adversarial input $\bm{\tilde{{I}}}$. $\operatorname{sign}(\cdot)$ is the sign indicator which has the value of 1 for positive value and -1 for negative value. The method creates subtle but impactful perturbations in autonomous driving.
Dot Attack[29] : Dot Attack targets deep learning algorithms by introducing minor, physical disruptions to camera systems. Specifically, a sticker with a designed dot pattern is placed on the camera lens.
Given an RGB image input $\bm{I}$, semi-transparent dots are strategically placed. We first consider the perturbation function for a single dot on the image $\bm{I}$. The perturbed image $\bm{\tilde{I}}$ at pixel location $(x_{i},x_{j})$ can be written as
$\begin{split}\bm{\tilde{I}}[x_{i},x_{j}]&=f_{0}(\bm{\phi_{0}};x_{i},x_{j})\\ &=(1-h(x_{i},x_{j}))\cdot\bm{I}(x_{i},x_{j})+h(x_{i},x_{j})\cdot\gamma\text{,}\end{split}$ (17)
where
$h(x_{i},x_{j})$ $=\alpha\cdot\exp\left(-\left(\frac{(x_{i}-x_{i}^{(c)})^{2}+(x_{j}-x_{j}^{(c)})^{2}}{r^{2}}\right)^{\beta}\right)$ (18)
$\bm{\phi_{0}}$ $=(\gamma,(x_{i}^{(c)},x_{j}^{(c)}),r,\alpha,\beta)\text{.}$ (19)
Here, $h(x_{i},x_{j})$ quantifies the balance between the original color at pixel $\bm{I}[x_{i},x_{j}]$ and a learnable color $\gamma$. This balance reduces with increasing distance from the center of the dot $\left(x_{i}^{(c)},x_{j}^{(c)}\right)$. $\bm{\phi_{0}}$ denotes the learnable parameters.
By composing $N_{\textrm{dot}}$ of these single-dot perturbations:
$f(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})=f(\bm{\phi_{0}};x_{i},x_{j})\circ...\circ f(\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{,}$ (20)
the pixel at the location $(x_{i},x_{j})$ of the final perturbed image $\bm{\tilde{I}}$ can be formalized as:
$\bm{\tilde{I}}[x_{i},x_{j}]=f_{0}(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{.}$ (21)
These parameters are learned by increasing the loss of the model, leading to poor performance of the trained model. In Fig. 2(b), the resulting dots mimic real-world lens stains, displaying a color gradient. Unlike computational attacks like FGSM, Dot Attacks are easily deployed in the real world, requiring only the application of a sticker to the camera lens.
V-A Experimental Design
In our experiments conducted within the CARLA simulation environment, we utilized a predefined trajectory, termed Town05 Long. To ensure a fair comparison across three distinct deep learning models—Transfuser+, Interfuser, and P-CSG—a consistent PID Controller was applied to all models. This uniform approach was designed to specifically assess the inherent robustness of each model under different adversarial attacks, by eliminating variations in control mechanisms.
FGSM Attacks: Our goal is to prompt immediate deviations in the vehicle’s trajectory by assigning target waypoints for the next four timesteps at $(-1.0,-2.0)$, $(-2.0,-4.0)$, $(-3.0,-6.0)$, and $(-4.0,-8.0)$, relative to the vehicle’s current position. These are used to formulate a dedicated loss function, from which adversarial perturbations are generated and applied to the input RGB images. We set the perturbation magnitude $\epsilon$ to 0.01 for these experiments.
Dot Attacks: Unlike FGSM Attacks, which necessitate per-frame computations, Dot Attacks employ a pre-trained, universally applicable Dot Mask. Our experimental procedure initiated with the nine-dots mask training phase, followed by its application within the CARLA simulation framework. This structured approach permitted us to evaluate the practicality and effectiveness of Dot Attack under conditions that closely mimic real-world scenarios.
V-B Results
In our evaluation, the Infraction Score serves as the primary metric, encapsulating various collision scenarios and traffic violations to comprehensively assess safety. In the context of adversarial attacks, a higher infraction score is crucial for the safety of both passengers and pedestrians. As shown in Table IV, our model, P-CSG, demonstrates superior performance in two distinct adversarial scenarios: high-intensity, per-frame FGSM Attack, and uniformly applied, low-intensity Dot Attack.
@@ V-B Results
Specifically, P-CSG achieved an Infraction Score that was at least $2.5$ times higher than that of competing models in both attack scenarios tested. Under the low-intensity Dot Attack, P-CSG not only maintained high safety standards, indicated by a higher Infraction Score but also completed more routes, as reflected in a higher Route Completion metric. Conversely, in scenarios involving high-intensity FGSM Attacks, P-CSG implemented a more conservative strategy aimed at enhancing security. This is shown by its higher Infraction Score, even though it led to fewer completed routes.
Most importantly, P-CSG consistently reaches the highest cumulative Driving Score in both types of attacks, confirming its robustness and adaptability in different challenging conditions.
VI Conclusion
@@
In this paper, we improve the fusion technologies and policy learning approaches based on penalties for autonomous driving. Additionally, we compare our proposed model with other baselines under Dot Attacks and FGSM Attacks, demonstrating outstanding performance and robustness. We aspire for our proposed penalty-based imitation learning approach to introduce a fresh perspective into the domain of end-to-end autonomous driving, aimed at enhancing the compliance of autonomous agents with traffic rules. Furthermore, additional research is necessary to enhance the robustness of autonomous driving systems, particularly in countering attacks such as Dot Attacks.

### v3

@@ 4.4 Ablation Studies
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$
No CSG $45.67\pm 5.51$ $84.82\pm 2.25$ $0.52\pm 0.10$
No Penalty $37.32\pm 3.58$ $78.41\pm 11.32$ $0.54\pm 0.21$
$\lambda_{1}=0.7$ $54.36\pm 1.78$ $84.37\pm 5.01$ $0.64\pm 0.04$
$\lambda_{1}=0.3$ $56.80\pm 2.22$ $89.60\pm 1.90$ $0.60\pm 0.02$
$\lambda_{2}=0.07$ $51.54\pm 1.50$ $81.49\pm 1.42$ $0.63\pm 0.02$
$\lambda_{2}=0.03$ $54.59\pm 2.88$ $88.12\pm 1.95$ $0.64\pm 0.02$
$\lambda_{3}=0.7$ $55.98\pm 3.00$ $\textbf{89.81}\pm 3.33$ $0.63\pm 0.02$
$\lambda_{3}=0.3$ $51.98\pm 1.93$ $84.66\pm 3.53$ $0.63\pm 0.01$
Table 3: Ablation Study.
two extra weights for each penalty are selected for comparison. We also provide the results of models without CSG and penalties for comprehensive analysis. The results of different penalty weights are also listed in the table. We found that assigning greater weight to more severe violations will increase the performance of our model. For instance, we apply greater penalties for the red light and the stop sign violations than overspeeding by turning since those two violations cause more serious consequences.
5 Robustness Study
Refer to caption(a)Refer to caption(b)Refer to caption(c)
Given the life-critical nature of autonomous driving, examining the robustness of neural network models against such vulnerabilities becomes essential. In this regard, we conduct a comprehensive comparative analysis to evaluate the robustness of our model. We benchmark its performance against two state-of-the-art methods — Transfuser+ and Interfuser. We focus on two specific types of white-box sensor attacks. These are Fast Gradient Sign Method (FGSM) Attack and Dot Attack.  [^1] Interfuser* has the same PID Controller as P-CSG and Transfuser+ to ensure a fair comparison.
  • FGSM Attack is notable for its computational efficiency and minimal perceptibility to the human eye. It uses the gradient of the model’s loss function to slightly alter the input image, creating an adversarial image that looks almost identical but can mislead the model.
  • Dot Attack uses special stickers with dot patterns on camera lenses to subtly blur parts of the image. These near-invisible dots mislead deep learning models without being noticeable to humans. Each dot’s color blends the original and a new RGB value, decreasing transparency from the center outward. This technique effectively disrupts model performance with minimal visual impact. Unlike computational attacks like FGSM, Dot Attacks are easily deployed in the real world, requiring only the application of a sticker to the camera lens.
5.1 Results
In our evaluation, the infraction score serves as the primary metric, encapsulating various collision scenarios and traffic violations to comprehensively assess safety. As shown in Table 4, P-CSG demonstrates superior performance in two distinct adversarial scenarios: high-intensity, per-frame FGSM Attack, and uniformly applied, low-intensity Dot Attack.
Attack Model Driving score Route compl. Infrac. score
@@ 5.1 Results
FGSM Attack $(\epsilon=0.01)$ P-CSG(Ours) $\mathbf{12.80}\pm 1.35$ $21.70\pm 2.87$ $\mathbf{0.59}\pm 0.02$
Transfuser+ $2.69\pm 0.73$ $42.44\pm 3.65$ $0.22\pm 0.06$
Interfuser* $8.37\pm 1.06$ $\mathbf{72.04}\pm 4.15$ $0.15\pm 0.04$
Dot Attack P-CSG(Ours) $\mathbf{36.60}\pm 3.17$ $\mathbf{75.53}\pm 6.43$ $\mathbf{0.56}\pm 0.07$
Transfuser+ $0.53\pm 0.21$ $54.06\pm 5.48$ $0.02\pm 0.01$
Interfuser* $12.56\pm 2.15$ $73.62\pm 7.32$ $0.19\pm 0.02$
Table 4: Attack Results.
@@
Specifically, P-CSG achieved an infraction score that was at least $2.5$ times higher than that of competing models in both attack scenarios tested. Under the low-intensity Dot Attack, P-CSG not only maintained high safety standards, indicated by a higher infraction score, but also completed more routes, as reflected in a higher route completion metric. Conversely, in scenarios involving high-intensity FGSM Attacks, P-CSG implemented a more conservative strategy aimed at enhancing security. Its higher infraction score shows this, even though it led to fewer completed routes. Most importantly, P-CSG consistently reaches the highest cumulative driving score in both types of attacks, confirming its robustness and adaptability in different challenging conditions.
6 Conclusion
@@
In this paper, we improve the multi-modality fusion technologies and policy learning methods based on penalties for autonomous driving. We observe that our contrastive learning-based multi-modal fusion method is helpful in extracting the global content of various modalities. Furthermore, we found that incorporating penalties based on traffic rules into an imitation learning pipeline can enhance the agent’s adherence to those traffic rules. We also compare our proposed model with other baselines under Dot Attacks and FGSM Attacks. The results show that our approach is more robust against these attacks. We aspire for our proposed penalty-based imitation learning approach to introduce a fresh perspective into the domain of end-to-end autonomous driving to enhance autonomous agents’ compliance with traffic rules.
One of our main objectives for future works is to include more types of penalties in the learning pipeline. These penalties could be designed based on multiple other traffic rules like speed limit, lane changing, and following distance. Scaling our method with a larger neural network and dataset is another direction of investigation, aimed at determining whether a larger model and data can effectively enhance overall performance.